quickstudent:

The difference of two numbers is 4. 5 times the smaller number is four more than three times the larger. Find the two numbers. -10 and -14 10 and 14 8 and 12 -8 and -12

Hero:

Let \(m\) represent the larger number and \(n\) represent the smaller number. \(m - n = 4\) \(5n = 3m + 4\) Sorry I have to remember different LaTeX formats for different sites. There's at least three different codes for rendering.

quickstudent:

Thanks! Now the answer I got is n = 8 and m = 12

Hero:

Mind showing how you arrived at that result?

Hero:

There's two ways to solve it. I just want to see which method you used.

quickstudent:

Sorry I took so long m - n = 4 m - n + n = 4 + n m = 4 + n 5n = 3(4 + n) + 4 5n = 12 + 3n + 4 5n = 16 + 3n 5n - 3n = 16 + 3n - 3n 2n = 16 2n/2 = 16/2 n = 8 m = 4 + 8 m = 12

Hero:

Very good. You used the correct method to find m and n which is called "Solving the system".

quickstudent:

Thanks!